# Randomly sampling bipartite networks with fixed degree sequences

### v4

@@ Abstract
journal: arXiv pre-print
@@ 2.1 Sampling bipartite networks using trade randomization
A network is bipartite when its nodes can be partitioned into two sets, such that edges exist only between nodes in different sets. A bipartite network is also two-mode when these two sets of nodes represent distinctly different kinds of entities. Such networks are common in the literature, and can capture many kinds of phenomena including event attendance where people are connected to events [Breiger, 1974], bill co-sponsorship where legislators are connected to bills [Neal, 2014], ecosystems where species are connected to habitats [Gotelli, 2000], and scholarship where authors are connected to papers [Newman, 2001]. Following Latapy et al. [2008], I generically refer to one set of nodes as ‘top nodes,’ and to the other set of nodes as ‘bottom nodes.’ A bipartite network is characterized by two degree sequences, one for the top nodes, and another for the bottom nodes. For example, in a bipartite network capturing event attendance, where people are top-nodes and events are bottom-nodes, the top-node degree sequence captures the number of events attended by each person, while the bottom-node degree sequence captures the number of people attending each event.
@@ 2.2 A trade randomization stopping rule
Typically $\mathcal{D}$ is unknown, so it is not possible to directly test whether the distribution of $D$ matches the distribution of $\mathcal{D}$. However, because $D$ converges on $\mathcal{D}$, it is only necessary to test when the distribution of $D$ has stabilized. There are many ways to test whether $D_{t}$ and $D_{t-1}$ were drawn from the same distribution. The Kolmogorov–Smirnov test [Kolmogorov, 1933, Smirnov, 1948] is the most widely known such test, however because lacks sensitivity in a range of cases, two other tests are more appropriate in this context. When the values in $D$ are unequally spaced (i.e., ordinal; as they are in the example illustrated in Figure 3), the Anderson-Darling (AD) test [Anderson & Darling, 1952] is optimal [Dowd, 2020]. In contrast, when the values in $D$ are equally spaced (i.e., interval), the DTS test is optimal [Dowd, 2020].
@@ 5 Discussion
Researchers performing statistical analysis of a bipartite network need to draw a random sample of all bipartite networks with the same degree sequences as the observed network. Efficient ‘trade’ algorithms exist for sampling bipartite networks through repeated randomization trades [Strona et al., 2014, Godard & Neal, 2022]. However, it is difficult to know how many trades needed to ensure the sample is random. In this paper, I proposed a stopping rule that focuses on the distribution of the sampled networks’ distances from a starting network, and stops performing trades once this distribution stabilizes. For a wide range of degree sequences, this stopping rule ensures with high probability (87.5% on average, 92% adjusted for type-I error) that the sample is a random sample of all bipartite networks with the given degree sequence. Additionally, in three larger empirical examples, the stopping rule performs as expected, and returns a sample in a practical amount of time.
The stopping rule offers a solution to a practical challenge encountered by researchers: determining when a sample of bipartite networks is random, regardless of how many trades this might require. These analyses do not directly answer a related question – how many trades are required, or what is the mixing time of a trade algorithm – but they do offer some insight. In their provided code, Strona et al. [2014] implicitly suggested that $5T$ trades was sufficient, where $T$ is the number of top nodes, but they did not provide any rationale or evidence for this suggestion. Was their suggestion reasonable?
For the small networks described in section 4.1, an average of about 4 trades per top node were required to obtain a random sample ($M=4.11,SD=0.50$). These results suggest that a $5T$ rule-of-thumb may be slightly liberal, but not unreasonable, in small networks. In contrast, for the empirical networks described in section 4.2, many more trades were required: $9T$ for Darwin’s Finches, $7T$ for Southern Women, and $11T$ for the US Senate. These results suggests that in larger networks a $5T$ rule-of-thumb may be too conservative, and that more trades are needed to achieve a random sample. While the results presented above do not provide definitive guidance on how many trades to perform in general, the stopping rule provides a way to determine how many trades to perform in a specific case, as well as a tool for further research on trade algorithm mixing times.
@@
Supplementary information, and all data and code necessary to replicate these results, are available at https://osf.io/rmwpz/?view_only=5d18c67044b44474addba3824cab2aa5.

### v5

# A stopping rule for randomly sampling bipartite networks with fixed degree sequences

@@ Abstract
This is a post-print. The version of record is available at: Neal, Z. P. (2024). A stopping rule for randomly sampling bipartite networks with fixed degree sequences. Social Networks, XX, XXX. https://doi.org/10.1016/j.socnet.XXX
@@
journal: Social Networks
@@ 2.1 Sampling bipartite networks using trade randomization
A network is bipartite when its nodes can be partitioned into two sets, such that edges exist only between nodes in different sets. A bipartite network is also two-mode when these two sets of nodes represent distinctly different kinds of entities. Such networks are common in the literature, and can capture many kinds of phenomena including event attendance where people are connected to events [Breiger, 1974], bill co-sponsorship where legislators are connected to bills [Neal, 2014], ecosystems where species are connected to habitats [Gotelli, 2000], and scholarship where authors are connected to papers [Newman, 2001]. Following Latapy et al. [2008], I generically refer to the smaller set of nodes as ‘top nodes,’ and to the larger set of nodes as ‘bottom nodes.’ A bipartite network is characterized by two degree sequences, one for the top nodes, and another for the bottom nodes. For example, in a bipartite network capturing event attendance, where people are top-nodes and events are bottom-nodes, the top-node degree sequence captures the number of events attended by each person, while the bottom-node degree sequence captures the number of people attending each event.
@@ 2.2 A trade randomization stopping rule
Typically $\mathcal{D}$ is unknown, so it is not possible to directly test whether the distribution of $D$ matches the distribution of $\mathcal{D}$. However, because $D$ converges on $\mathcal{D}$, it is only necessary to test when the distribution of $D$ has stabilized. There are many ways to test whether $D_{t}$ and $D_{t-1}$ were drawn from the same distribution. The Kolmogorov–Smirnov test [Kolmogorov, 1933, Smirnov, 1948] is the most widely known such test. However, because this test lacks sensitivity in a range of cases, two other tests are more appropriate in this context. When the values in $D$ are unequally spaced (i.e., ordinal; as they are in the example illustrated in Figure 3), the Anderson-Darling (AD) test [Anderson & Darling, 1952] is optimal [Dowd, 2020]. In contrast, when the values in $D$ are equally spaced (i.e., interval), the DTS test is optimal [Dowd, 2020].
@@ 5 Discussion
Researchers performing statistical analysis of a bipartite network need to draw a random sample of all bipartite networks with the same degree sequences as the observed network. Efficient ‘trade’ algorithms exist for sampling bipartite networks through repeated randomization trades [Strona et al., 2014, Godard & Neal, 2022]. Although using these algorithms on contemporary computers can perform trades quite fast, and future algorithms executed using future computing technology will enable trades to be performed even faster, it is still necessary to know how many trades needed to ensure the sample is random. In this paper, I proposed a stopping rule that focuses on the distribution of the sampled networks’ distances from a starting network, and stops performing trades once this distribution stabilizes. For a wide range of degree sequences, this stopping rule ensures with high probability (87.5% on average, 92% adjusted for type-I error) that the sample is a random sample of all bipartite networks with the given degree sequence. Additionally, in three larger empirical examples, the stopping rule performs as expected, and returns a sample in a practical amount of time.
The stopping rule offers a solution to a practical challenge encountered by researchers: determining when a sample of bipartite networks is random, regardless of how many trades this might require. Importantly, this challenge cannot be solved by greater computational power alone. Greater computational power may allow trades to be performed more quickly, thereby enabling faster sampling and sampling of larger bipartite networks, but researchers still require a rule for deciding how many trades to perform.
These analyses do not directly answer a related question – how many trades are required, or what is the mixing time of a trade algorithm – but they do offer some insight. In their provided code, Strona et al. [2014] implicitly suggested that $5T$ trades was sufficient, where $T$ is the number of top nodes, but they did not provide a rationale or evidence for this suggestion. Was their suggestion reasonable?
For the small networks described in section 4.1, an average of about 4 trades per top node were required to obtain a random sample ($M=4.11,SD=0.50$). These results suggest that a $5T$ rule-of-thumb may be slightly liberal, but not unreasonable, in small networks. In contrast, for the empirical networks described in section 4.2, many more trades were required: $9T$ for Darwin’s Finches, $7T$ for Southern Women, and $11T$ for the US Senate. These results suggests that in larger networks a $5T$ rule-of-thumb may be too conservative, and that more trades are needed to achieve a random sample. More generally, contrary to the suggestion from Strona et al. [2014], these results suggest that there is not a linear relationship between the number of top nodes and the number of trades required. While the results presented above do not provide definitive guidance on how many trades to perform, the stopping rule provides a way to determine how many trades to perform in a specific case, as well as a tool for further research on trade algorithm mixing times.
@@
Supplementary information, and all data and code necessary to replicate these results, are available at https://osf.io/rmwpz/.